### v4

Abstract
Deepfake is a technology dedicated to creating highly realistic facial images and videos under specific conditions, which has significant application potential in fields such as entertainment, movie production, and digital human creation, to name a few. With the advancements in deep learning, techniques primarily represented by Variational Autoencoders and Generative Adversarial Networks have achieved impressive generation results. More recently, the emergence of diffusion models with powerful generation capabilities has sparked a renewed wave of research. In addition to deepfake generation, corresponding detection technologies continuously evolve to regulate the potential misuse of deepfakes, such as privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing current state-of-the-arts in this rapidly evolving field. First, we unify task definitions, comprehensively introduce datasets and metrics, and discuss developing technologies. Then, we discuss the development of several related sub-fields and focus on researching four representative deepfake fields: face swapping, face reenactment, talking face generation, and facial attribute editing, as well as forgery detection. Subsequently, we comprehensively benchmark representative methods on popular datasets for each field, fully evaluating the latest and most influential published works. Finally, we analyze the challenges and future research directions of the discussed fields. We closely follow the latest developments in this project.
Keywords:
Deepfake Generation Face Swapping Face Reenactment Talking Face Generation Facial Attribute Editing Forgery detection Survey
1 Introduction
Artificial Intelligence Generated Content (AIGC) garners considerable attention sha2023deep in academia and industry. Deepfake generation, as one of the important technologies in the generative domain, gains significant attention due to its ability to create highly realistic facial media content. This technique transitions from traditional graphics-based methods to deep learning-based approaches. Early methods employ advanced Variational Autoencoder kingma2014auto ; sohn2015learning ; van2017neural (VAE) and Generative Adversarial Network (GAN) karras2019style ; karras2020analyzing techniques, enabling seemingly realistic image generation, but their performance is still unsatisfactory, which limits practical applications. Recently, the diffusion structure blattmann2023stable ; guo2023animatediff ; liu2024towards has greatly enhanced the generation capability of images and videos. Benefiting from this new wave of research, deepfake technology demonstrates potential value for practical applications and can generate content indistinguishable from real ones, which has further attracted attention and is widely applied in numerous fields barattin2023attribute , including entertainment, movie production, online live broadcasting, and privacy protection, etc.
Figure 1: Time diagram that reflects the survey pipeline. Zoom in for a better holistic perception of this work.
Deepfake generation can generally be divided into four mainstream research fields: 1) Face swapping xu2022region ; ancilotto2023ximswap ; shiohara2023blendface is dedicated to executing identity exchanges between two person images; 2) Face reenactment koujan2020head2head ; hsu2022dual ; bounareli2023hyperreenact emphasizes transferring source movements and poses; 3) Talking face generation mir2023dit ; ma2023dreamtalk ; zhang2024dr2 focuses on achieving natural matching of mouth movements to textual content in character generation, and 4) Facial attribute editing sun2022anyface ; huang2023ia ; pang2023dpe aims to modify specific facial attributes of the target image. The development of related foundational technologies has gradually shifted from single forward GAN models goodfellow2014generative ; karras2019style to multi-step diffusion models ho2020denoising ; rombach2022high ; blattmann2023stable with higher quality generation capabilities, and the generated content has also gradually transitioned from single-frame images to temporal video modeling guo2023animatediff . In addition, NeRF mildenhall2020nerf ; gao2022nerf has been frequently incorporated into modeling to improve multi-view consistency capabilities jiang2022nerffaceediting ; zhang2022fdnerf .
While enjoying the novelty and convenience of this technology, the unethical use of it raises concerns over the spread of privacy invasion, and dissemination of fake news, necessitating the development of effective forgery detection methods tan2023learning in opposition, a.k.a., deepfake detection. From the earliest handcrafted feature-based methods he2019detection ; zhou2017two to deep learning-based methods yin2023dynamic ; yang2023masked , and the recent hybrid detection techniques ilyas2023avfakenet , forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains qian2020thinking ; li2021frequency to the more challenging temporal domain haliassos2022leveraging ; yang2023avoid . Considering that current generative technologies have a higher level of interest, develop faster, and can generate indistinguishable content from reality mirsky2021creation , corresponding detection technologies need continuous evolution.
Overall, despite the significant progress made in both directions, they still exhibit challenging limitations in specific scenarios vyas2024analysing , mainly reflected in the visual perception authenticity and generative accuracy of models. This has attracted a large number of researchers to continue their efforts and has sparked thoughts on industrial applications. Existing survey works only focus on partial deepfake fields and lack discussions on new technologies sha2023deep ; mirsky2021creation ; liu2023gan , especially diffusion-based image/video generation methods, due to their disconnection from current technologies. The survey will comprehensively discuss these fields as well as related sub-fields, and cover tracking of the latest works.
Figure 2: Top: Illustration of different deepfake generation (Sec. 3.1) and detection tasks (Sec. 3.2) that are discussed in this survey. Bottom: Specific facial attribute modification of each task. Data from NVIDIA Keynote at COMPUTEX 2023 at 29:40.
$\bullet$ Contribution. In this survey, we comprehensively explore the key technologies and latest advancements in Deepfakes generation and forgery detection. We first unify the task definitions (Sec. 2.1), provide a comprehensive comparison of datasets and metrics (Sec. 2.3), and discuss the development of related technologies. Specifically, we investigate four mainstream deepfake fields: face swapping (Sec. 3.1.1), face reenactment (Sec. 3.1.2), talking face generation (Sec. 3.1.3), and facial attribute editing (mainly on multiple editing) (Sec. 3.1.4), as well as forgery detection (Sec. 3.2). We also analyze the benchmarks and settings for each domain, thoroughly evaluating the latest and influential works published in top-tier conferences/journals (Sec. 4), especially recent diffusion-based approaches. Additionally, we discuss closely related fields, including head swapping, face super-resolution, face reconstruction, face inpainting, body animation, portrait style transfer, makeup transfer, and adversarial sample detection. Benefiting from the current popularity of AIGC, the research iteration cycle in the deepfake field has been significantly reduced, and we keep updating and discussing the in-submission works in the revised version.
$\bullet$ Scope. This survey primarily focuses on mainstream face-related tasks, including face swapping, face reenactment, talking face generation, facial (multiple) attribute editing, and forgery detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et alsha2023deep only discuss character generation while we cover a more comprehensive range of tasks. Compared to works mirsky2021creation ; liu2023gan ; melnik2024face , our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Pipeline. Fig. 1 shows the pipeline of this survey. Sec. 2 provides the essential background knowledge that encompasses task-specific definitions, datasets, evaluation metrics, and exploration of related research areas. Sec. 3 presents technical discussions, examining the four most popular deepfake tasks from the perspective of technological categorization and evolution. In addition, we meticulously classify and discuss forgery detection techniques, emphasizing the technological routes involved. Subsequently, Sec. 4 organizes and evaluates performance from various approaches for a thorough and fair performance comparison. Sec. 5 focuses on a critical review of the challenges that persist in existing techniques, outlining feasible future development directions. Finally, we encapsulate the entirety of the paper through a comprehensive summary in Sec. 6.
2 Background
In this section, we first introduce the conceptual definitions of the discussed mainstream fields. Fig. 2 illustrates the intuitive objectives for each task and shows the distinctions among tasks in terms of manipulated facial components. Then, we review the developmental history of commonly used neural networks, highlighting several representative ones. Next, we summarize popular datasets, metrics, and loss functions. Finally, we comprehensively discuss several relevant domains.
2.1 Problem Definition
$\bullet$ Unified Formulation of Studied Problems. Fig. 2 intuitively displays the various deepfakes generation and detection tasks studied in this paper. For the former, different tasks can essentially be expressed as controlled content generation problems under specific conditions, such as images, audio, text, specific attributes, etc.. Given the target image $I_{t}$ to be manipulated and the condition information $C=\{\mathtt{Image},\mathtt{Audio},\mathtt{Text},\dots\}$, the content generation process can be represented as:
$I_{o}={\bm{\phi_{G}}}(I_{t},C),$ (1)
where $\phi_{G}$ abstracts the specific generation network and $I_{o}=\{I_{t}^{0},I_{t}^{1},\dots,I_{t}^{N-1}\}$ represents generated contents. $N$ is the total frame number for the generated video, which is set to 1 by default). The latter task can be viewed as an image-level or pixel-level classification problem as practical application needs, which can be represented as:
$S_{o}={\bm{\phi_{D}}}(I_{o}),$ (2)
where $\phi_{D}$ abstracts the specific detection network and $S_{o}$ represents the fake score for the generated content $I_{o}$.
$\bullet$ Face Swapping involves replacing the identity information of the target face $I_{t}$ by the source face $C$=$I_{s}$, while preserving ID-irrelevant attributes of the target face, such as skin color and expressions.
$\bullet$ Face Reenactment alters the facial movement of a target image $I_{t}$ without changing the identity and other facial attributes, under the conditions of driving image, video, head pose, etc.. This technology often involves the support of facial motion capture technology, such as facial tracking or prediction based on deep learning models zheng2021efficient .
Figure 3: Development timeline of three mainstream generative modals, i.e., VAE, GAN, and Diffusion.
$\bullet$ Talking Face Generation can be viewed as an extension in time, aiming at generating a talking video $I_{o}$=$\{I_{o}^{i}\},i=0,1,\cdots,N-1$ with the character in the target image $I_{t}$ engaging in a conversation based on an arbitrary driving source, such as text, audio, video, or a multi-modal composite source. The lip movements, facial poses, expressions, emotions, and spoken content information of the character in the generated video match the target information.
$\bullet$ Facial Attribute Editing aims to modify the semantic information of the target face $I_{t}$ (e.g., personality, age, expressions, skin color, etc.) in a directed manner based on individual interest and preference. Existing methods include single and comprehensive attribute editing: the former focuses on training a model for only one attribute, while the latter integrates multiple attribute editing tasks simultaneously that is our primary focus in this survey.
$\bullet$ Forgery Detection aims to detect and identify anomalies, tampering, or forgery areas in images or videos by the anomaly score $S_{o}$, and it has great research and application value in information security and multimedia forensics.
2.2 Technological History and Roadmap
$\bullet$ Generative Framework. Variational Autoencoders (VAEs) kingma2014auto ; sohn2015learning ; van2017neural , Generative Adversarial Networks (GANs) goodfellow2014generative ; karras2019style ; karras2020analyzing , and Diffusion sohl2015deep ; ho2020denoising ; rombach2022high have played pivotal roles in the developmental history of generative models.
1) VAE kingma2014auto emerges in 2013, altering the relationship between latent features and linear mappings in autoencoder latent spaces. It introduces feature distributions like the Gaussian distribution and achieves the generation of new entities through interpolation. To enhance generation capabilities under specific conditions, CVAE sohn2015learning introduces conditional input. VQ-VAE van2017neural introduces the concept of vector quantization to improve the learning of latent representations. Subsequent models continually advance.
2) GANs goodfellow2014generative achieve high-quality generation through adversarial training with an extra discriminator. Subsequently, research on GANs experiences a surge, and currently, implementations based on GANs remain the mainstream approach for various deepfake tasks. CGAN mirza2014conditional introduces conditional control variables to GANs. Pix2Pix isola2017image enhances GAN performance in specific image translation tasks. StyleGAN karras2019style introduces the concept of style transfer to GANs. StyleGAN2 karras2020analyzing further improves the quality and controllability of generated images. Additionally, some composite research combines GANs with VAEs, e.g., CVAE-GAN bao2017cvae .
3) Diffusion models sohl2015deep model the generation of data as a diffusion process. DDPM ho2020denoising gains widespread attention for its outstanding generative performance, especially when the model excels in handling large-scale, high-resolution images. LDM rombach2022high is more flexible and powerful when it comes to modeling complex data distributions. In the field of video generation, diffusion models play a crucial role ma2023follow ; liu2023residual ; gupta2023photorealistic . SVD blattmann2023stable fine-tunes the base model using an image-to-video conversion task based on text-to-video models. AnimateDiff guo2023animatediff attaches a newly initialized motion modeling module to a frozen text-to-image model and trains it on subsequent video clips to refine reasonable motion prior knowledge, thus achieving excellent generation results.
Figure 4: Works summarization on different directions per year. Data is obtained on 2024/05/13.
Table 1: Overview of commonly used datasets. Orange-marked ones are selected to evaluate different methods in Sec. 4.
Dataset Type Scale Highlight
Deepfake Generation LFW huang2008labeled Image 10K Facial images captured under various lighting conditions, poses, expressions, and occlusions at 250$\times$250 resolution.
CelebA liu2015deep Image 200K The dataset includes over 200K facial images from more than 10K individuals, each with 40 attribute labels.
VGGFace parkhi2015deep Image 2600K A super-large-scale facial dataset involving a staggering 26K participants, encompassing a total of 2600K facial images.
VoxCeleb1 nagrani2017voxceleb Video 100K voices A large scale audio-visual dataset of human speech, the audio includes noisy background interference.
CelebA-HQ karras2018progressive Image 30K A high-resolution facial dataset consisting of 30K face images, each with a resolution of 1024$\times$1024 resolution.
VGGFace2 cao2018vggface2 Image 3000K A large-scale facial dataset, expanded with greater diversity in terms of ethnicity and pose compared to VGGFace.
VoxCeleb2 chung2018voxceleb2 Video 2K hours A dataset that is five times larger in scale than VoxCeleb1 and improves racial diversity.
FFHQ karras2019style Image 70K The dataset with over 70K high-resolution (1024$\times$1024) facial images, showcasing diverse ethnicity, age, and backgrounds.
MEAD wang2020mead Video 40 hours An emotional audiovisual dataset provides facial expression information during conversations with various emotional labels.
CelebV-HQ zhu2022celebv Video 35K videos The dataset’s video clips have a resolution of more than 512$\times$512 resolution and are annotated with rich attribute labels.
Forgery Detection DeepfakeTIMIT korshunov2018deepfakes Audio-Video 640 videos The dataset is evenly divided into two versions: LQ (64$\times$64) and HQ (128$\times$128), with all videos using face swapping forgery.
FF++ rossler2019faceforensics++ Video 6K videos Comprising 1K original videos and manipulated videos generated using five different forgery methods.
DFDCP dolhansky2019deepfake Audio-Video 5K videos The preliminary dataset for The Deepfake Detection Challenge includes two face-swapping methods.
DFD DfD Video 3K videos The dataset comprises 3K deepfake videos generated using five forgery methods.
Deeperforensics jiang2020deeperforensics Video 60K videos The videos feature faces with diverse skin tones, and rich environmental diversity was considered during the filming process.
DFDC dolhansky2020deepfake Audio-Video 128K videos The official dataset for The Deepfake Detection Challenge and contain a substantial amount of interference information.
Celeb-DF li2020celeb Video 1K videos Comprising 408 genuine videos from diverse age groups, ethnicities, and genders, along with 795 DeepFake videos.
Celeb-DFv2 li2020celeb Video 6K videos An expanded version of Celeb-DFv1, this dataset not only increases in quantity but also the diversity.
FakeAVCeleb khalid2021fakeavceleb Audio-Video 20K videos A novel audio-visual multimodal deepfake detection dataset, deepfake videos generated using four forgery methods.
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) lecun1998gradient ; krizhevsky2012imagenet ; he2016deep ; liu2022convnet ; emo ; eatformer have played a pivotal role in the history of deep learning. LeNet lecun1998gradient , as the pioneer of CNNs, showcased the charm of machine learning. AlexNet krizhevsky2012imagenet and ResNet he2016deep made deep learning feasible. Recently, ConvNeXt liu2022convnet has achieved excellent results surpassing those of Swin-Transformer liu2021swin . The Transformer architecture initially proposed vaswani2017attention in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT dosovitskiy2020image demonstrates that using Transformer in the field of computer vision can still achieve excellent performance and PVT wang2021pyramid overcomes the challenges of adapting Transformer to various dense prediction tasks. In addition, Swin-Transformer liu2021swin addresses the limitations of Transformer in handling high-resolution image processing tasks. Subsequently, Swin-Transformer V2 liu2022swin further improves the model’s efficiency and the resolution of manageable inputs.
$\bullet$ Neural Radiance Field (NeRF) is first introduced in 2020 mildenhall2020nerf , with its core idea revolving around the use of volume rendering and implicit neural fields to represent and reconstruct both geometric and illumination information of 3D scenes gao2022nerf . Compared to traditional 3D methods, it exhibits higher visual quality and is currently widely applied in tasks such as 3D geometry enhancement sharma2023neural ; irshad2023neo , segmentation liu2023instance and 6D pose estimation min2023neurocs . In addition, Some notable works hwang2023faceclipnerf ; yang2023urbangiraffe combining NeRF as a supplement to 3D information and generation models are particularly prominent at present.
$\bullet$ Work Summary. The evolution of mainstream generative models is depicted chronologically in Fig. 3. This survey delves into four categories of generation tasks along with the forgery detection task, and the publication years distribution of the surveyed articles is shown in Fig. 4.
2.3 Datasets, Metrics, and Losses
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW huang2008labeled , CelebA liu2015deep , CelebA-HQ karras2018progressive , VGGFace parkhi2015deep , VGGFace2 cao2018vggface2 , FFHQ karras2019style , Multi-PIE moore2010multi , VoxCeleb1 nagrani2017voxceleb , VoxCeleb2 chung2018voxceleb2 , MEAD wang2020mead , MM CelebA-HQ xia2021tedigan , CelebAText-HQ sun2021multi , CelebV-HQ zhu2022celebv , TalkingHead-1KH wang2021one , LRS2 afouras2018deep , LRS3 afouras2018lrs3 , etc. 2) Commonly used forgery detection datasets include UADFV li2018ictu , DeepfakeTIMIT korshunov2018deepfakes , FF++ rossler2019faceforensics++ , Deeperforensics-1.0 jiang2020deeperforensics , DFDCP dolhansky2019deepfake , DFDC dolhansky2020deepfake , Celeb-DF li2020celeb , Celeb-DFv2 li2020celeb , FakeAVCeleb khalid2021fakeavceleb , DFD DfD , WildDeepfake zi2020wilddeepfake , KoDF kwon2021kodf , UADFV yang2019exposing , Deephy narayan2022deephy , DF-Platter narayan2023df , etc. We summarize popular datasets in Table 1 .
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) wang2023memory , Structured Similarity (SSIM) wang2004image , Learned Perceptual Image Patch Similarity (LPIPS) zhang2018unreasonable , Fréchet Inception Distance (FID) heusel2017gans , Kernel Inception Distance (KID) binkowski2018demystifying , Cosine Similarity (CSIM) wang2021one , Identity Retrieval Rate (ID Ret) wang2018cosface , Expression Error chaudhuri2019joint , Pose Error ruiz2018fine , Landmark Distance (LMD) around the mouths chen2018lip , Lip-sync Confidence (LSE-C) prajwal2020lip , Lip-sync Distance (LSE-D) prajwal2020lip , etc. 2) For forgery generation commonly uses: Area Under the ROC Curve (AUC) le2023quality , Accuracy (ACC) wang2022m2tr , Equal Error Rate (EER) huang2023implicit , Average Precision (AP) tan2023learning , F1-Score chen2021image , etc. Detailed definitions are explained in Sec. 4.1.
$\bullet$ Loss Function. VAE-based approaches generally employ reconstruction loss and KL divergence loss kingma2014auto . Commonly used reconstruction loss functions include Mean Squared Error, Cross-Entropy, LPIPS zhang2018unreasonable , and perceptual johnson2016perceptual losses. GAN-based methods further introduce adversarial loss goodfellow2014generative to increase image authenticity, while diffusion-based works introduce denoising loss function ho2020denoising .
2.4 Related Research Domains
$\bullet$ Head Swapping replaces the entire head information rather than only the face of the target image shu2022few , including facial contours and hairstyle, with that of the source image. However, in terms of facial attributes, Head Swap only replaces identity attributes, while leaving other attribute information unchanged. Recently, some methods wang2023hs ; han2023generalist based on Diffusion have been proposed and have shown promising results in achieving good performance.
$\bullet$ Face Super-resolution aims to enhance the resolution of low-resolution face images to generate high-resolution face images scsnet ; moser2023hitchhiker . This task is closely related to various deepfake sub-tasks. In the early stages of Face Swapping natsume2018rsgan ; sun2018hybrid and Talking Face Generation fan2015photo ; chen2019hierarchical , there is consistently a low-resolution issue in the generated images and videos. This problem is addressed by incorporating face super-resolution methods into the models mir2023dit ; xu2022high to improve output quality. In terms of technical approaches, FSR can be categorized into methods based on CNNs huang2017wavelet ; chen2019sequential ; chen2020learning , GANs KoMulti ; moser2024diffusion , reinforcement learning shi2019face , and ensemble learning jiang2019atmfn .
$\bullet$ Face Reconstruction refers to the process of recreating the three-dimensional appearance of an individual’s face based on one or multiple 2D facial images morales2021survey ; sharma20223d . Facial reconstruction often plays an intermediate role in various deepfake sub-tasks. In Face Swapping and Face Reenactment tasks, facial parameters are reconstructed using 3DMM, and the model’s parameters can be controlled directionally. Additionally, reconstructing a 3D facial model is one of the methods to address the issue of facial artifacts in videos generated under large facial poses. Common technical approaches for facial reconstruction include methods based on 3DMM maninchedda2017fast ; lin2020towards , epipolar geometry feng20183d , one-shot learning tulsiani2017multi ; xing2019self , shadow shape reconstruction kemelmacher20103d ; jiang20183d , and hybrid learning-based reconstruction dou2017end ; chaudhuri2020personalized .
$\bullet$ Face Inpainting, a.k.a. face completion, aims to reconstruct missing regions in face images caused by external factors such as occlusion and lighting while preserving facial texture information is crucial in this process zhang2023image . This task is a crucial sub-task of image inpainting, and the current methods are mostly based on deep learning that can be roughly divided into two categories: GAN based yildirim2023diverse ; zhou2023superior and Diffusion based xu2024personalized ; yang2024pgdiff .
$\bullet$ Body Animation aims to alter the entire bodily pose while unchanging the overall body information yu2023bidirectionally . The goal is to achieve a modification of the target image’s entire body posture using an optional driving image or video, aligning the body of the target image with the information from the driving signal. The mainstream implementation path for body animation is based on GANs zhou2022cross ; zhao2022thin , and Diffusion wang2023leo ; ma2023follow ; wang2023disco ; xu2023magicanimate .
$\bullet$ Portrait Style Transfer aims to reconstruct the style of a target image to match that of a source image by learning the stylistic features of the source image yang2022vtoonify ; peng2023portraitbooth . The goal is to preserve the content information of the target image while adapting its style to that of the source image cai2023image . Common applications include image cross-domain style transfer, such as transforming real face images into animated face styles perez2024styleavatar ; feng20243d . Methods based on GANs abdal20233davatargan ; xu2022your3demoji ; jiang2023scenimefy and Diffusion liu2023portrait ; hur2024expanding have achieved high-quality performance in this task.
$\bullet$ Makeup Transfer aims to achieve style transfer learning from a source image to a target image jiang2020psgan ; li2023hybrid . Existing models have achieved initial success in applying and removing makeup on target images chang2018pairedcyclegan ; gu2019ladn ; liu2021psgan++ ; li2023hybrid , allowing for quantitative control over the intensity of makeup. However, they perform poorly in transferring extreme styles zhong2023sara ; yang2022elegant ; yan2023beautyrec . Existing mainstream methods are based on GANs sun2023ssat ; hao2022cumtgan ; jiang2020psgan .
$\bullet$ Adversarial Sample Detection focuses on identifying whether the input data is an adversarial sample han2023interpreting . If recognized as such, the model can refuse to provide services for it, such as throwing an error or not producing an output yuan2021current . Current deepfake detection models often rely on a single cue from the generation process as the basis for detection, making them vulnerable to specific adversarial samples. Furthermore, relatively little work has focused on adversarial sample testing in terms of model generalization capability and detection evaluation.
3 Deepfake Inspections: A Survey
This section systematically examines four generation tasks: Face Swapping (Sec. 3.1.1), Talking Face Generation (Sec. 3.1.3), Face Reenactment (Sec. 3.1.2), Facial Attribute Editing (Sec. 3.1.4). Additionally, we review techniques for forgery detection (commonly known as deepfake detection) in Sec. 3.2. We document and compare the most representative works in the main text. Furthermore, we discuss several related domains that garnered significant attention in Sec. 3.3.
3.1 Deepfake Generation
Table 2: Overview of representative face swapping methods. Notations: ➊ Self-build, ➋ CelebA-HQ, ➌ FFHQ, ➍ VGGFace2, ➎ VGGFace, ➏ CelebV, ➐ CelebA, ➑ VoxCeleb2, ➒ LFW, ➓ KoDF. Abbreviations: SIGGRAPH (SIG.), EUROGRAPHICS (EG.), GANs (G.), VAEs (V.), Diffusion (D.), Split-up and Integration (SI.).
Method Venue Dataset Categorize Limitation Highlight
Traditional Graphics Blanz et alblanz2004exchanging EG.’04 3DMM Manual intervention, unnatural output. Early face-swapping efforts simplified manual interaction steps.
Bitouk et albitouk2008face SIG.’08 SI. Manual intervention, attribute loss. A three-phase implementation framework with the help of a pre-constructed face database to match faces that are similar to the source face in terms of posture and lighting.
Sunkavalli et alsunkavalli2010multi TOG’10 SI. Poor generalizability, frequent artifacts. Early work on face exchange was realized using image processing methods such as smooth histogram matching technique.
Dale et aldale2011video SIG.’11 3DMM Poor generalizability and output quality. Early work on face exchange, proposing an improved Poisson mixing approach to achieve face swapping in video through frame-by-frame face replacement.
Lin et allin2012face ICME’12 zhu2009unsupervised 3DMM Poor generalizability, frequent artifacts. An attempt to construct a personalized 3D head model to solve the artifact problem occurring in face swapping in large poses.
Mosaddegh et almosaddegh2015photorealistic ACCV’14 gross2010multi milborrow2010muct SI. Poor generalizability and output quality. A diverse form of face swapping where facial components can be targeted for replacement.
Nirkin et alnirkin2018face FG’18 burgos2013robust SI. Poor generalization ability and resolution. Transfer of expressions and poses by building some 3D variable models and training facial segmentation networks to maintain target facial occlusion.
Generative Adversarial Network IPGAN bao2018towards CVPR’18 guo2016ms G.+V. Poor output image quality, frequent artifacts. Using two encoders to encode facial identity and attribute information separately for facial information decoupling and swapping.
RSGAN natsume2018rsgan SIG.’18 G.+V. Loss of lighting information. Using two independent VAE modules to represent the latent spaces of the face and hair regions, respecti -vely, with the replacement of identity information in the latent space implemented.
Sun et alsun2018hybrid ECCV’18 zhang2015beyond G.+3DMM Poor ability to preserve face feature attributes. Implementing in two stages: the first stage involves replacing the identity information of the face reg -ion, while the second stage achieves complete facial rendering.
FSGAN nirkin2019fsgan ICCV’19 maze2018iarpa G. Poor ability to preserve face feature attributes. Two novel loss functions are introduced to refine the stitching in the face fusion phase following the swapping process.
FaceShifter li2019faceshifter CVPR’20 ➊➋➌➎ G. Poor ability to preserve face feature attributes. Face swapping is realized in two stages, the firstly AEI-Net improves the output image quality level, and the second HEAR-Net is targeted to focus on abnormal regions for image recovery.
Zhu et alzhu2020deepfakes AAAI’20 G.+V. Inability to process facial contour information. First show of the applicability of deepfake to keypoint invariant de-identification work.
DeepFaceLab liu2023deepfacelab PR’20 G.+V. Poor output image resolution. A mature and complete framework for face swapping.
SimSwap chen2020simswap MM’20 ➊➍ G.+V. Poor ability to preserve face feature attributes. ID modules and weak feature matching loss functions are proposed to find a balance between identity information replacement and attribute information retention.
MegaFS zhu2021one CVPR’21 ➊➋➌ G. Poor ability to preserve face feature attributes. The first method allows for face swapping on images with a resolution of one million pixels.
FaceInpainter li2021faceinpainter CVPR’21 ➊➋➌➎ G.+3DMM Poor representation of image details. A two-stage framework innovatively implements heterogeneous domains face swapping.
HifiFace wang2021hififace IJCAI’21 G.+3DMM Uses a large number of parameters. A 3D shape-aware identity extractor is proposed to achieve better retention of attribute information such as facial shape.
FSGANv2 nirkin2022fsganv2 TPAMI’22 maze2018iarpa G. Unable to process posture differences effectively. An extension of the FSGAN method that combines Poisson optimization with perceptual loss enhances the output image facial details.
StyleSwap xu2022styleswap ECCV’22 ➊➎➑ G. Unable to process posture differences effectively. Proposing an exchange-guided ID reversal strategy to enhance the performance of attribute informa -tion replacement during face exchange.
RAFSwap xu2022region CVPR’22 G. Unable to process posture differences effectively. Local facial region awareness branch with global source feature adaptation (SFA) branch is proposed to better achieve the preservation of target image attribute information.
FSLSD xu2022high CVPR’22 ➊➋ G. Poor ability to preserve face feature attributes. Potential semantic de-entanglement is realized to obtain facial structural attributes and appearance attributes in a hierarchical manner.
Kim et alkim2022smooth CVPR’22 ➌➍ G. Unable to process posture differences effectively. An identity embedder is proposed to enhance the training speed under supervision.
3DSwap li20233d CVPR’23 ➋➌ G.+3DMM Unable to process posture differences effectively. A 3d-aware approach to the face-swapping task, de-entangling identity and attribute features in latent space to achieve identity replacement and attribute feature retention.
FALCO barattin2023attribute CVPR’23 ➋➒ G. Poor ability to handle facial occlusion. Oriented with privacy-preserving applications, the method directly employs latent space of pre-trained GANs to achieve the identity of anonymized images while preserving facial attributes.
WSC-Swap ren2023reinforced ICCV’23 ➊➋➌➎ G.+3DMM Poor resolution of the output image . Two mutually independent encoders are proposed to encode attribute information outside the face region and semantic-level non-identical facial attributes inside the face region.
BlendFace shiohara2023blendface ICCV’23 ➊➌➍➏ G. Unable to handle occlusion and extreme lighting. The identity features obtained from the de-entanglement are fed to the generator as an identity loss function, which guides the generator to generate an image to fit the source image identity information.
FlowFace zeng2023flowface AAAI’23 ➊➋➌➍ G.+3DMM Altered target image lighting details. It consists of face reshaping network and face exchange network, which better solves the influence of the difference between source and target face contours on the face exchange work.
S2Swap liu2023high MM’23 ➊➋➌➑ G.+3D Poor ability to preserve face feature attributes. Achieving high-fidelity face swapping through semantic disentanglement and structural enhancement.
StableSwap zhu2024stableswap TMM’24 ➊➌ G.+3D Unable to handle extreme skin color differences. Utilizing a multi-stage identity injection mechanism effectively combines facial features from both the source and target to produce high-fidelity face swapping.
Difussion DiffFace kim2022diffface arXiv’22 ➊➌ D. Facial lighting attributes are altered. Claims to be the first diffusion model-based face exchange framework.
DiffSwap zhao2023diffswap CVPR’23 ➊➌ D. Poor ability to handle facial occlusion. Reenvisioning face swapping as conditional inpainting to harness the power of the diffusion model.
FaceX han2023generalist arXiv’23 ➊➌➏ D. Unable to handle extreme skin color differences. A novel facial all-rounder model capable of performing various facial tasks.
Liu et alliu2024towards arXiv’24 ➊➋➌ D. Poor ability to preserve face feature attributes. Conditional diffusion model introduces identity and expression encoders components, achieving a balan- ce between identity replacement and attribute preservation during the generation process.
Alternative Cui et alcui2023face CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware corresponden- ces between source and target faces.
TransFS cao2023transfs FG’23 ➊➋➓ Other Unable to process posture differences effectively. The identity generator is designed to reconstruct high-resolution images of specific identities, and an attention mechanism is utilized to enhance the retention of identity information.
Wang et alwang2024efficient TMM’24 ➊➋ Other Poor ability to handle facial occlusion. A Global Residual Attribute-Preserving Encoder (GRAPE) is proposed, and a network flow considering the facial landmarks of the target face was introduced, achieving high-quality face swapping.
3.1.1 Face Swapping
In this section, we review face swapping methods from the perspective of basic architecture, which can be mainly divided into four categories and summarized in  Table 2.
$\bullet$ Traditional Graphics. As representative early implementations, traditional graphics methods in the implementation path can be divided into two categories: 1) Key information matching and fusion. Methods bitouk2008face ; sunkavalli2010multi ; nirkin2018face ground in critical information matching and fusion are geared towards substituting corresponding regions by aligning key points within facial regions of interest (ROIs), such as the mouth, eyes, nose, and mouth, between the source and target images. Following this, additional procedures such as boundary blending and lighting adjustments are executed to produce the resulting image. Bitouk et albitouk2008face accomplish automated face replacement by constructing a substantial face database to locate faces with akin poses and lighting conditions for substitution. Meanwhile, Nirkin et alnirkin2018face enhance keypoint matching and segmentation accuracy by incorporating a Fully Convolutional Network (FCN) into their method. 2) The construction of a 3D prior model for facial parameterization. Methods blanz2004exchanging ; dale2011video based on constructing a 3D prior and introducing a facial parameter model often involve building a facial parameter model using 3DMM technology based on a pre-collected face database. After matching the facial information of the source image with the constructed face model, specific modifications are made to the relevant parameters of the facial parameter model to generate a completely new face. Dale et aldale2011video utilize 3DMM to track facial expressions in two videos, enabling face swapping in videos. Some methods lin2012face ; guo2019face explore scenarios involving significant pose differences between the source and target images. Lin et allin2012face construct a 3D face model from frontal faces, renderable in any pose. Guo et alguo2019face utilize plane parameterization and affine transformation to establish a one-to-one dense mapping between 2D graphics. Traditional computer graphics methods solve basic face-swapping problems, exploring full automation to enhance generalization. However, these methods are constrained by the need for similarities in pose and lighting between source and target images. They also face challenges like low image resolution, modification of target attributes, and poor performance in extreme lighting and occlusion scenarios.
$\bullet$ Generative Adversarial Network. GAN-based methods aim to obtain realistic images generated through adversarial training between the generator and the discriminator, becoming the mainstream face swapping approach. According to different improvement objectives, methods can be classified into seven categories:
1) Early GAN-based methods bao2017cvae ; moniz2018unsupervised ; liu2023deepfacelab ; zhu2020deepfakes address issues related to the similarity of pose and lighting between source and target images. DepthNets moniz2018unsupervised combines GANs with 3DMM to map the source face to any target geometry, not limited to the geometric shape of the target template. This allows it to be less affected by differences in pose between the source and target faces. However, they face challenges in generalizing the trained model to unknown faces.
2) Improved Generalizability. To improve the model’s generalization, many efforts sun2018hybrid ; nirkin2019fsgan ; chen2020simswap are made to explore solutions. Combining GANs with VAEs, the model bao2018towards ; natsume2018rsgan encodes and processes different facial regions separately. FSGAN nirkin2019fsgan integrates face reenactment with face swap, designing a facial blending network to mix two faces seamlessly. SimSwap chen2020simswap introduces an identity injection module to avoid integrating identity information into the decoder. However, these methods suffer from low resolution and significant attribute loss and need help to handle facial occlusions effectively.
3) Resolution Upgrading. Some methods zhu2021one ; xu2022designing ; jiang2023styleipsb provide solutions to enhance the resolution of generated images. MegaFS zhu2021one introduces the first single-lens face swapping method at the million-pixel level. The face encoder no longer compresses facial information but represents it in layers, achieving more detailed preservation. StyleIPSB jiang2023styleipsb constrains semantic attribute codes within the subspace of StyleGAN, thereby fixing certain semantic information during face swapping to preserve pore-level details.
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods wang2021hififace ; ren2023reinforced ; zeng2023flowface ; zhang2023flowface++ introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace wang2021hififace introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace zeng2023flowface introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ zhang2023flowface++ improves upon FlowFace by utilizing a pre-trained Mask Autoencoder to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
5) Facial Masking Artifacts. Some methods li2019faceshifter ; nirkin2019fsgan ; liu2023fine ; rosberg2023facedancer have partially alleviated the artifacts caused by facial occlusion. FSGAN nirkin2019fsgan designs a restoration network to estimate missing pixels. E4S liu2023fine redefines the face-swapping problem as a mask-exchanging problem for specific information. It utilizes a mask-guided injection module to perform face swapping in the latent space of StyleGAN. However, overall, the methods above have not thoroughly addressed the issue of artifacts in generated images under extreme occlusion conditions.
6) Trade-offs between Identity-replacement and Attribute-retention. In addition to the occlusion issues that need further handling, researchers gao2021information ; shiohara2023blendface discover that the balance between identity replacement and attribute preservation in generated images seems akin to a seesaw. Many methods xu2022styleswap ; kim2022smooth ; liu2023high explore the equilibrium between identity replacement and attribute retention. InfoSwap gao2021information aims to decouple identity and attribute information in face swapping by leveraging the information bottleneck theory. It seeks controlled swapping of identity between source and target faces. StyleSwap xu2022styleswap introduces a novel swapping guidance strategy, the ID reversal, to enhance the similarity of facial identity in the output. Shiohara et alshiohara2023blendface propose BlendFace, using an identity encoder that extracts identity features from the source image and uses it as identity distance loss, guiding the generator to produce facial exchange results.
7) Model Light-weighting is also an important topic with profound implications for the widespread application of models. FastSwap yoo2023fastswap achieves this by innovating a decoder block called Triple Adaptive Normalization (TAN), effectively integrating identity information from the source image and pose information from the target image. XimSwap ancilotto2023ximswap modifies the design of convolutional blocks and the identity injection mechanism, successfully deploying on STM32H743.
Table 3: Overview of representative face reenactment methods. Notations: ➊ Voxceleb, ➋ Self-build, ➌ Voxceleb2, ➍ TalkingHead-1KH, ➎ CelebV-HQ, ➏ VFHQ, ➐ RaFD, ➑ VGGFace, ➒ CelebV, ➓FFHQ. Abbreviations: Expression (Exp).
Method Venue Controllable object Dataset Highlight
Based on 3DMM
Face2Face thies2016face2face CVPR’16 Exp, Pose Using 3D facial reconstruction, transfer facial expressions and pose from a driving character to a target character via affine transforma -tion, then generate the final video through rendering techniques.
Kim et alkim2018deep TOG’18 Exp, Pose, Blink Using synthesized rendering images of a parameterized face model as input, creating lifelike video frames for the target actor.
Kim et alkim2019neural TOG’19 Lip,Exp, Pose Built on a recurrent generative adversarial network, it employs a hierarchical neural face renderer to synthesize realistic video frames.
Head2Head koujan2020head2head FG’21 Exp, Pose, Blink, Gaze The model consists of two stages: facial reconstruction and tracking in the first stage, followed by video rendering in the second stage.
HeadGAN doukas2021headgan ECCV’21 Exp, Pose Using 3DMM for facial modeling provides a 3D prior to the GAN, effectively guiding the generator to accurately recover pose and expression from the target frame.
Face2Faceρ yang2022face2face ECCV’22 Exp, Pose Decoupling the actor’s facial appearance and motion information with two separate encodings allows the network to learn facial app -earance and motion priors.
PECHead gao2023high CVPR’23 Lip, Exp, Pose ➌➍➎➏ A novel multi-scale feature alignment module for motion perception is proposed to minimize distortion during motion transmission.
Based on Landmark Matching
X2Face wiles2018x2face CVPR’18 Exp, Pose The training is realized in two stages: the first stage guides the generation frames towards the driving frames and the second stage accomplishes the preservation of the source identity information through the identity loss function.
Zakharov et alzakharov2019few ICCV’19 Exp, Pose ➊➌ Proposed a meta-learning framework for adversarial generative models, reducing the required training data size.
FReeNet zhang2020freenet CVPR’20 Exp, Pose moore2010multi A new triple perceptual loss is proposed to richly reproduce facial details of the face.
Zakharov et alzakharov2020fast ECCV’20 Exp, Pose Decomposed facial information into two layers for modeling: the first layer synthesizes coarse images related to the pose using a small neural network, and the second layer defines texture images unrelated to the pose, containing high-frequency details.
MarioNETte ha2020marionette AAAI’20 Exp, Pose ➊➒ Introduced image attention blocks, target feature alignment modules, and landmark transformers, enhancing the model’s performance when generalizing to unknown individual identities.
DG hsu2022dual CVPR’22 Exp, Pose ➊➌➐ A proposed dual generator model network for large pose face reproduction.
Doukas et aldoukas2023free TPAMI’23 Exp, Pose, Gaze kellnhofer2019gaze360 zhang2015appearance Eye gaze control in the generated video is implemented to further enhance visual realism.
MetaPortrait zhang2023metaportrait CVPR’23 Exp, Pose By establishing dense facial keypoint matching, accurate deformation field prediction is achieved, and the model training is expedited based on the meta-learning philosophy.
Yang et alyang2023learning AAAI’24 Exp, Pose ➊➌➍ The facial tri-plane is represented by canonical tri-plane, identity deformation, and motion components, achieving face reenactment without the need for 3D parameter model priors.
FSRTrochow2024fsrt CVPR’24 Exp, Pose The Transformer-based encoder-decoder effectively encodes attributes and improves action transmission quality.
Based on Face Feature Decopling
HiDe-NeRF li2023one CVPR’23 Lip, Exp, Pose ➊➌➍ High-fidelity and free-viewing talking head synthesis using deformable neural radiation fields.
HyperReenact bounareli2023hyperreenact ICCV’23 Exp, Pose ➊➌ Exploiting the effectiveness of hypernetworks in real image inversion tasks and extending them to real image manipulation.
Stylemask bounareli2023stylemask FG’23 Exp, Pose This work optimizes a Mask Network and combines it with StyleGAN2’s style potential space S in order to achieve the separation of facial pose and expression of the target image from the identity features of the source image.
Bounareli et albounareli2024one IJCV’24 Exp, Pose ➊➌ In GANs’ latent space, head pose and expression changes are decoupled, achieving near-real outputs through real image embedding.
Based on Self-supervised
ICface tripathy2020icface WACV’20 Exp, Pose The model is decoupled and driven by interpretable control signals that can be obtained from multiple sources such as external driving videos and manual controls.
Oorloff et aloorloff2023robust ICCV’23 Lip, Exp, Pose Identity and attribute decomposition are realized in StyleGAN2’s latent space, and a cyclic manifold adjustment technique enhances facial reconstruction results.
Xue et al. xue2023high TOMM’23 Exp, Pose ➊➌ High-fidelity facial generation is achieved by using information-rich Projected Normalized Coordinate Code (PNCC) and eye maps, replacing sparse facial landmark representations.
$\bullet$ Diffusion-based. The latest studies kim2022diffface ; zhao2023diffswap ; han2023generalist ; liu2024towards in this area produce promising generation results. DiffSwaps zhao2023diffswap redefines the face swapping problem as a conditional inpainting task. Liu et alliu2024towards introduce a multi-modal face generation framework and achieved this by introducing components such as balanced identity and expression encoders to the conditional diffusion model, striking a balance between identity replacement and attribute preservation during the generation process. As a novel facial generalist model, FaceX han2023generalist can achieve various facial tasks, including face swapping and editing. Leveraging the pre-trained StableDiffusion blattmann2023stable has significantly improved the quality and model training speed.
$\bullet$ Alternative Techniques. Some methods stand independently from the above classifications that are discussed here collectively. Fast Face-swap korshunova2017fast views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods cui2023face ; cao2023transfs apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer liu2021swin , TransFS cao2023transfs obtains rich facial features, enabling facial swapping in high-resolution images. ReliableSwap yuan2023reliableswap enhances the model’s identity preservation capabilities by constructing a reliable supervisor called the "cyclic triplet." However, it has limitations in preserving attribute information.
3.1.2 Face Reenactment
This section reviews current methods from four points: 3DMM-based, landmark matching, face feature decoupling, and self-supervised learning. We summarize them in Table 3.
$\bullet$ 3DMM-based. Some methods koujan2020head2head ; yang2022face2face ; gao2023high utilize 3DMM to construct a facial parameter model as an intermediary for transferring information between the source and target. In particular, Face2Faceρ yang2022face2face , based on 3DMM, consists of a u-shaped rendering network driven by head pose and facial motion fields and a hierarchical coarse-to-fine motion network guided by landmarks at different scales. However, some methods thies2016face2face ; kim2018deep ; kim2019neural exhibit visible artifacts in the background when dealing with significant head movements in the input images. To address issues such as incomplete attribute decoupling in facial reproduction tasks, PECHead gao2023high models facial expressions and pose movements. It combines self-supervised learning of landmarks with 3D facial landmarks and introduces a new motion-aware multi-scale feature alignment module to eliminate artifacts that may arise from facial motion.
$\bullet$ Landmark Matching. This kind of methods wiles2018x2face ; zakharov2020fast ; ha2020marionette ; hsu2022dual ; doukas2023free aim to establish a mapping relationship between semantic objects in the facial regions of the driving source and the target source through landmarks. Based on this mapping relationship zakharov2019few ; zhang2020freenet , the transfer of facial movement information is achieved. In particular, X2Face wiles2018x2face achieves one-to-one driving of facial expressions and poses for each frame of the character in the driving video and source video. To address the challenge of reproducing large head poses in facial reenactment, Xu et alhsu2022dual propose a dual-generator network incorporating a 3D landmark detector into the model. Free-headgan doukas2023free comprises a 3D keypoint estimator, an eye gaze estimator, and a generator built on the HeadGAN architecture. The 3D keypoint estimator addresses the regression of deformations related to 3D poses and expressions. The eye gaze estimator controls eye movement in videos, providing finer details. MetaPortrait zhang2023metaportrait achieves accurate distortion field prediction through dense facial keypoint matching and accelerates model training based on meta-learning principles, delivering excellent results on limited datasets.
$\bullet$ Feature Decoupling. The latent feature decoupling and driving methods bounareli2023hyperreenact ; bounareli2023stylemask ; li2023one ; bounareli2024one ; yang2023learning aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact bounareli2023hyperreenact uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask bounareli2023stylemask separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF li2023one employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
$\bullet$ Self-supervised Learning. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels tripathy2020icface ; zeng2020realistic ; oorloff2023robust ; zhang2023exploiting . Oorloff et aloorloff2023robust employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et alzhang2023exploiting utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide motion field estimation, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
Table 4: Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ.
Method Venue Dataset Limitation Highlight
Audio / Text - Driven Chen et alchen2018lip ECCV’18 cooke2006audio 0Audiovisual Poor resolution, inability to control pose and emotional. Proposed a novel generator network and a comprehensive model with four complementary losses, as well as a new audio-visual related loss function to guide video generation.
Zhou et alzhou2019talking AAAI’19 Inability to control pose and emotional variations. Generate high-quality talking face videos by disentangling audio-visual representations.
Chen et alchen2019hierarchical CVPR’19 Inability to control pose and emotional variations. Proposed a cascaded approach, using facial landmarks as an intermediate high-level representation.
Wav2Lip prajwal2020lip ICMR’20 ➊➎➐ Poor resolution, inability to control pose and emotional. A new evaluation framework and a dataset for training mouth synchronization are proposed.
MakeItTalk zhou2020makelttalk TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
Ji et alji2021audio CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
SPACE gururani2023space ICCV’23 ➋➌ Lack emotional and other latent attributes control Constructed a novel facial intermediate representation, achieving control overhead pose, blinking, and gaze direction.
Gan et algan2023efficient ICCV’23 ➊➋➌ Insufficient emotional output diversity. Proposed a two-stage architecture, implementing speaker generation independent of emotion and emb -edding emotion information.
SadTalker zhang2023sadtalker CVPR’23 ➏➒ Lack emotional and other latent attributes control. Based on the idea of 3DMM and conditional VAE, 3D coefficients controlling facial motion and expre -ssion are generated from audio to realize the reproduction of accurate faces from audio.
Zhongetczhong2023identity CVPR’23 ➎➐ afouras2018deep Lack emotional and other latent attributes control. A two-stage framework for step-by-step audio-to-face action generation is proposed.
EmoTalk peng2023emotalk ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
TalkLip wang2023seeing CVPR’23 ➊➎ Inability to control pose and emotional variations. Pre-trained lip-reading experts are employed to penalize incorrect lip-reading predictions in the synth -esized videos.
DR2 zhang2024dr2 WACV’24 Lack emotional and other latent attributes control. The model explored effective strategies for reducing the training workload.
RADIO lee2024radio WACV’24 ➊➋➏ Lack emotional and other latent attributes control. Introducing StyleGAN2 style modulation to adapt to human identity and utilizes ViT blocks to focus on facial attributes in the reference image.
Multimodal PC-AVS zhou2021pose CVPR’21 ➊➋ Lack emotional and other latent attributes control. Introduction of pose-source video drive compensation to generate head motion in video.
GC-AVT liang2022expressive CVPR’22 ➋➌ Poor resolution, unable to handl complex backgrounds. In addition to the source image, a gesture source, an expression source, and audio are introduced to drive the talking head generation.
Yu et alyu2021multimodal TMM’22 Lack emotional and other latent attributes control. Fusion of audio and text inputs for more accurate lip movement and chin posture prediction.
Xu et alxu2023high CVPR’23 Insufficient control over the intensity of emotional output. Embedding textual, visual, and auditory emotional modalities into a unified space.
LipFormer wang2023lipformer CVPR’23 ➎➓ Poor ability to preserve face feature attributes. Propose retaining high-quality facial details obtained from pre-training in a codebook format and repro -ducing them by driving the encoded mapping relationship between audio and lip movements.
Wang et alwang2024styletalk++ TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
Diffusion DAE-Talker du2023dae MM’23 High model complexity. It replaces traditional manually crafted intermediate representations with data-driven latent representa -tions obtained from a DAE.
Yu et alyu2023talking ICCV’23 ➋➒ Poor resolution, high model complexity. Building a corresponding mapping between audio and non-lip representations and training using the diffusion model.
DreamTalk ma2023dreamtalk arXiv’23 ➋➌➏ Mismatched emotion and semantics, occasional artifacts. The denoising network, style-aware lip expert, and style predictor collaborate to make the model perform well in various speaking styles.
Stypułkowski stypulkowski2024diffused WACV’24 ➊➑ High model complexity, short video generation duration. The model incorporates motion frame and audio embedding information to capture past movements and future expressions, with an emphasis on the mouth region through an additional lip sync loss.
EmoTalker zhang2024emotalker ICASSP’24 ➌➑ High model complexity. It achieves emotion-editable talking face generation based on a conditional diffusion model.
VASA-1 xu2024vasa arXiv’24 High model complexity. Expressive and well-decoupled facial latent space has been constructed, and highly controllable, high -quality generation effects have been achieved based on the Diffusion Transformer.
3D-Model AD-NeRF guo2021ad ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
DFRF shen2022learning ECCV’22 Lack of emotional and other latent attributes control. Combining audio with 3D perceptual features and proposing an facial deformation module.
AE-NeRF li2023ae AAAI’24 Lack emotional and other latent attributes control. Facial modeling is divided into NeRF related to audio and unrelated to audio to enhance audio-visual lip synchronization and facial detail.
SyncTalk peng2023synctalk CVPR’24 Lack controllable emotional intensity regulation. The facial sync controller boosts component coordination, and a portrait generator corrects artifacts, enhancing video details.
Ye et alye2024real3d ICLR’24 zhu2022celebv Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.
3.1.3 Talking Face Generation
In this section, we review current methods from three perspectives: audio/text driven, multimodal conditioned, diffusion-based, and 3D-model Technologies. We also summarize them in Table 4.
$\bullet$ Audio/Text Driven. Methods aim to map and guide lip and facial movements in generated videos by understanding the semantic information from the driving source apb2face ; shen2022learning ; wang2023seeing . Early methods fan2015photo ; chen2018lip perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers chen2019hierarchical ; prajwal2020lip propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk zhou2020makelttalk decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. SadTalker zhang2023sadtalker extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods wu2023speech2lip ; wang2023memory ; lee2024radio ; fu2023mimic propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos sheng2023stochastic ; gan2023efficient . Recently, some methods tan2023emmn ; zhai2023talking ; gururani2023space extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN tan2023emmn establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO zhai2023talking employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP ma2023talkclip introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et alzhong2023identity propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 zhang2024dr2 explores practical strategies for reducing the training workload.
$\bullet$ Multimodal Conditioned. To generate more realistic talking videos, some methods zhou2021pose ; liang2022expressive ; xu2023high ; wang2023lipformer introduce additional modal information on top of audio-driven methods to guide facial pose and expression. GC-AVT liang2022expressive generates realistic talking videos by independently controlling head pose, audio information, and facial expressions. This approach introduces an expression source video, providing emotional information during the speech and the pose source video. However, the video quality falls below expectations, and it struggles to handle complex background changes. Xu et alxu2023high integrate text, image, and audio-emotional modalities into a unified space to complement emotional content in textual information. Multimodal approaches have significantly enhanced the vividness of generated videos, but there is still room for exploration of organically combining information driven by different sources and modalities.
$\bullet$ Diffusion-based. Recently, some methods mir2023dit ; ma2023dreamtalk ; du2023dae ; xu2023multimodal ; zhang2024emotalker apply the Diffusion model to the task of talking face generation. For fine-grained talking video generation, DAE-Talker du2023dae replaces manually crafted intermediate representations, such as facial landmarks and 3DMM coefficients, with data-driven latent representations obtained from a Diffusion Autoencoder (DAE). The image decoder generates video frames based on predicted latent variables. EmoTalker zhang2024emotalker utilizes a conditional diffusion model for emotion-editable talking face generation. It introduces emotion intensity blocks and the FED dataset to enhance the model’s understanding of complex emotions. Very recently, diffusion models are gaining prominence in talking face generation tasks stypulkowski2024diffused ; wei2024aniportrait ; tian2024emo and video generation tasks wang2023leo ; karras2023dreampose . Emo tian2024emo directly predicts video from audio without the need for intermediate 3D components, achieving excellent results. However, the lack of explicit control signals may easily lead to unnecessary artifacts. Based on the Diffusion Transformer architecture, VASA-1 xu2024vasa finely encodes and reconstructs facial details, constructing an expressive and well-decoupled facial latent space. This results in highly controllable and high-resolution video generation.
$\bullet$ 3D-model Technologies. 3D models, exemplified by NeRF, are gaining traction in talking face generation guo2021ad ; shen2022learning ; li2023ae . AD-NeRF guo2021ad directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic NeRF. AE-NeRF li2023ae employs a dual NeRF framework to separately model audio-related regions and audio-independent regions. Furthermore, some methods peng2023synctalk ; ye2024real3d adopt Tri-Plane chan2022efficient to represent facial and audio attributes. Synctalk peng2023synctalk models and renders head motion using a tri-plane hash representation, and then further enhances the output quality using a portrait synchronization generator. Very recently, 3D Gaussian Splatting kerbl20233d also been widely applied to this task. Some method chen2024gstalker ; cho2024gaussiantalker ; li2024talkinggaussian ; yu2024gaussiantalker introduce 3DGS to achieve more refined facial reconstruction and motion details, aiming to address the issue of insufficient pose and expression control caused by NeRF’s implicit representation.
Table 5: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.).
Method Venue Categorize Dataset Highlight
SC-FEGAN jo2019sc ICCV’19 G. Users can generate high-quality edited output images by freely sketching parts of the source image.
AttGAN he2019attgan TIP’19 G. Applying attribute classification constraints to generated images has validated the drawbacks of enforcing stringent attribute independence constraints in latent representations.
Shen et alshen2020interpreting CVPR’20 G. Thoroughly investigated how to encode different semantics in the latent space and explored the disentanglement between various semantics to achieve precise control over facial attributes.
Yao et alyao2021latent ICCV’21 G.+T. By integrating explicit decoupling terms and identity-consistent terms into the loss function, the preservation of facial identity information is improved, resulting in high-quality face editing in videos.
HifaFace gao2021high CVPR’21 G. Proposed a solution based on wavelet transform to address the issue of partial loss of attribute information when generating edited results due to "cyclic consistency" problems.
Preechakul et alpreechakul2022diffusion CVPR’22 D. When encoding images, it is divided into semantically meaningful parts and parts that represent the details of the image.
FENeRF sun2022fenerf CVPR’22 G.+NeRF ➊➍ The introduction of semantic masks into the conditional radiance field enables finer image textures.
GuidedStyle hou2022guidedstyle NN’22 G. Generating faces after face editing is guided based on facial attribute classification. The introduction of a sparse attention mecha -nism enhances the manipulation of individual attribute styles.
FDNeRF zhang2022fdnerf SIG.’22 G.+NeRF The introduction of the Conditional Feature Warping (CFW) module addresses the issue of temporal inconsistency caused by dynamic information in the process of face editing in videos.
AnyFace sun2022anyface CVPR’22 G. ➏➑ Proposed a dual-branch framework for text-driven facial editing, with coordination achieved between the two branches through a Cross-Modal Distillation (CMD) module.
TransEditor xu2022transeditor CVPR’22 G. ➌➊ Emphasizing dual-space GAN interaction’s importance, a transformer architecture is introduced for improved interaction.
Huang et alhuang2023collaborative CVPR’23 D. Proposed the concept of assisted diffusion, integrating individual multimodalities to explore the complementarity between differe -nt modalities.
Ozkan et alozkan2023conceptual ICCV’23 G. The entangled attribute space is decomposed into conceptual and hierarchical latent spaces, and transformer network encoders are employed to modify information in the latent space.
CIPS-3D++ zhou2023cips TPAMI’23 G.+NeRF ➊➒ Replaced the convolutional architecture with an MLP (Multi-Layer Perceptron) architecture to achieve faster rendering speeds.
ClipFace aneja2023clipface SIG.’23 G.+3DMM Learned texture generation from large-scale datasets, enhancing generator performance through generative adversarial training.
TG-3DFace yu2023towards ICCV’23 G. ➌➏ For different scenarios, two sets of text-to-face cross-modal alignment methods were designed with specific focuses.
VecGAN++ dalva2022vecgan TPAMI’23 G. Orthogonal constraint and disentanglement loss are used to decouple attribute vectors in the latent space.
DiffusionRig ding2023diffusionrig CVPR’23 D. 3DMM and diffusion model integration propose a two-stage method for learning personalized facial details.
Kim et alkim2023diffusion CVPR’23 D. Proposed a method for facial editing in videos based on the diffusion model.
SDGAN et alHuang2024SDGAN AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
FaceDNeRF zhang2024facednerf NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.
3.1.4 Facial Attribute Editing
In this section, we review current methods chronologically, following the progression in overcoming technical challenges, primarily focusing on multiple attribute editing methods utilizing GANs. Finally, we summarize methods in Table 5.
$\bullet$ Comprehensive Editing. Facial attribute editing aims to selectively alter specific facial attributes without affecting others. Therefore, disentangling different facial attributes is a primary challenge. Early facial attribute editing models shen2017learning ; zhou2017genegan often achieve editing for a single attribute through data-driven training. For instance, Shen et alshen2017learning propose learning the difference between pre-/post-operation images, represented as residual images, to achieve attribute-specific operations. However, single-attribute editing falls short of meeting expectations, and compression steps in the process often lead to a significant loss of image resolution, a common issue in early methods. The fundamental challenge in comprehensive editing and unrelated attribute modification is achieving complete attribute disentanglement. Many approaches shen2020interpreting ; yao2021latent ; gao2021high ; xu2022transeditor have explored this. E.g., HifaFace gao2021high identifies cycle consistency issues as the cause of facial attribute information loss that proposes a wavelet-based method for high-fidelity face editing, while TransEditor xu2022transeditor introduces a dual-space GAN structure based on the transformer framework that improves image quality and attribute editing flexibility.
$\bullet$ Irrelevant-attribute Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images yao2021latent ; hou2022guidedstyle ; huang2023ia . GuidedStyle hou2022guidedstyle leverages attention mechanisms in StyleGAN karras2019style for the adaptive selection of style modifications for different image layers. IA-FaceS huang2023ia embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches sun2022fenerf ; zhang2022fdnerf ; jiang2022nerffaceediting ; zhou2023cips combine GANs with NeRF mildenhall2020nerf for enhanced spatial awareness capabilities. Specifically, FENeRF sun2022fenerf uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ zhou2023cips enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
$\bullet$ Text Driven facial attribute editing is a crucial application scenario and a recent hot topic in academic research sun2022anyface ; hou2022textface ; aneja2023clipface ; yu2023towards . TextFace hou2022textface introduces text-to-style mapping, directly encoding text descriptions into the latent space of pre-trained StyleGAN. TG-3DFace yu2023towards introduces two text-to-face cross-modal alignment techniques, including global contrastive learning and fine-grained alignment modules, to enhance the high semantic consistency between the generated 3D face and the input text.
$\bullet$ Diffusion-based models have been introduced into facial attribute editing preechakul2022diffusion ; huang2023collaborative ; ding2023diffusionrig ; kim2023diffusion and achieve excellent results. Huang et alhuang2023collaborative propose a collaborative diffusion framework, utilizing multiple pre-trained unimodal diffusion models together for multimodal face generation and editing. DiffusionRig ding2023diffusionrig conditions the initial 3D face model, which helps preserve facial identity information during personalized editing of facial appearance based on the general facial details before the dataset.
Table 6: Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD.
Method Venue Train Test Highlight
Space Domain Gram-Net liu2020global CVPR’20 ➆➈ ➆➈ The method posits that genuine faces and fake faces exhibit inconsistencies in texture details.
Face X-ray li2020face CVPR’20 ➀➁➂➉ Focusing on boundary artifacts of face fusion for forgery detection.
Zhao et alzhao2021multi CVPR’21 ➀➁➂ A texture enhancement module, an attention generation module, and a bilinear attention pooling mod -ule are proposed to focus on texture details.
Nirkin et alnirkin2021deepfake TPAMI’21 ➀➁➂ Detecting swapped faces by comparing the facial region with its context (non-facial area).
SBIs shiohara2022detecting CVPR’22 ➁➂➇➉ The belief that the more difficult to detect forged faces typically contain more generalized traces of forg -ery can encourage the model to learn a feature representation with greater generalization ability.
RECCE cao2022end CVPR’22 ➀➁➂zi2020wilddeepfake Reconstruction learning on real samples to learn common compressed representations of real images.
LGrad tan2023learning CVPR’23 The gradient is utilized to present generalized artifacts that are fed into the classifier to determine the truth of the image.
NoiseDF wang2023noise AAAI’23 ➀➁➂➃ Extracting noise traces and features from cropped faces and background squares in video frames.
Ba et alba2024exposing AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on information bottleneck theory, is proposed for constraint.
Time Domain Yang et alyang2019exposing ICASSP’19 Focusing on the inconsistency in the head pose in videos by comparing the estimated head pose using all facial landmarks with the one estimated using only the landmarks in the central region.
FTCN zheng2021exploring ICCV’21 ➀➁➂➃li2019faceshifter It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
LipForensics haliassos2021lips CVPR’21 ➀➁➂ Concern about temporal inconsistency of mouth movements in videos.
M2TR wang2022m2tr ICMR’22 ➀➁➂➉ Capturing local inconsistencies at different scales for forgery detection using a multiscale transformer.
Gu et algu2022delving AAAI’22 ➀➁➂zi2020wilddeepfake By densely sampling adjacent frames to pay attention to the inter-frame image inconsistency.
Yang et alyang2023masked TIFS’23 ➀➁➂ ➀➁➂ Treating detection as a graph classification problem and focusing on the relationship between the local image features across different frames.
AVoiD-DF yang2023avoid TIFS’23 ➁➄khalid2021fakeavceleb ➁➄khalid2021fakeavceleb Multimodal forgery detection using audiovisual inconsistency.
Choi et alchoi2024exploiting CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
Xu et alxu2024towards IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
Peng et alpeng2024deepfakes TIFS’24 ➀➂➇ ➀➂➇ Focuse on inter-frame gaze angles, extracting gaze informations and employing spatio-temporal feature aggregation to combine temporal, spatial, and texture features for detection and classification.
Frequency F3-Net qian2020thinking ECCV’20 A two-branch frequency perception framework with a cross-attention module is proposed.
FDFL li2021frequency CVPR’21 Propose an adaptive frequency feature generation module to extract differential features from different frequency bands in a learnable manner.
HFI-Net miao2022hierarchical TIFS’22 ➁➂➃➅korshunov2018deepfakes Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
Guo et alguo2023constructing TIFS’23 ➀➁ ➀➁➂ Designing a backbone network for Deepfake detection with space-frequency interaction convolution.
Tan et altan2024frequency AAAI’24 ➀➄ A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.
Data Driven Dang et aldang2020detection CVPR’20 ➂➅ Utilizing attention mechanisms to handle the feature maps of the detection model.
Zhao et alzhao2021learning ICCV’21 ➀➁➂➃➇➉ Proposes pairwise self-consistent learning for training CNN to extract these source features and detect deep vacation images.
Finfer hu2022finfer AAAI’22 ➀➂➇zi2020wilddeepfake Based on an autoregressive model, using the facial representation of the current frame to predict the facial representation of future frames.
Huang et alhuang2023implicit CVPR’23 ➀➁➂➉li2019faceshifter A new implicit identity-driven face exchange detection framework is proposed.
HiFi-Net guo2023hierarchical CVPR’23 Converting forgery detection and localization into a hierarchical fine-grained classification problem.
Zhai et alzhai2023towards ICCV’23 dong2013casia hsu2006detecting wen2016coverage Weakly supervised image processing detection is proposed such that only binary image level labels (real or tampered) are required for training.
3.2 Forgery Detection
In this section, we review current forgery detection techniques based on the type of detection cues, categorizing them into three: Space Domain (Sec. 3.2.1), Time Domain (Sec. 3.2.2), Frequency Domain (Sec. 3.2.3), and Data-Driven (Sec. 3.2.4). We also summarize the detailed information about popular methods in Table 6.
3.2.1 Space Domain
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color he2019detection , saturation mccloskey2019detecting , artifacts zhao2021multi ; shiohara2022detecting ; cao2022end , gradient variations tan2023learning , etc. Specifically, RECCE cao2022end considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad tan2023learning utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions nirkin2021deepfake , as well as the fine-grained details of image textures chai2020makes ; liu2020global . Recently, Ba et alba2024exposing focuse not only on the discordance in a single image region but also on the detection of fused local representation information from multiple non-overlapping areas.
$\bullet$ Local Noise Inconsistency. Image forgery may involve adding, modifying, or removing content in the image, potentially altering the noise distribution in the image. Detection methods based on noise aim to identify such local or even global differences in the image. Zhou et alzhou2017two propose a dual-stream structure, combining GoogleNet with a triplet network to focus on tampering artifacts and local noise in images. Nguyen et alnguyen2019capsule utilize capsule networks to detect forged pictures and videos in various forgery scenarios. NoiseDF wang2023noise specializes in identifying underlying noise traces left behind in Deepfake videos, introducing an efficient and novel Multi-Head Relative Interaction with depth-wise separable convolutions to enhance detection performance.
3.2.2 Time Domain
$\bullet$ Abnormal Physiological Information. Forgery videos often overlook the authentic physiological features of humans, failing to achieve overall consistency with authentic individuals. Therefore, some methods focus on assessing the plausibility of the physiological features of the generated faces in videos. Li et alli2018ictu detect blinking and blink frequency in videos as criteria for determining the video’s authenticity. Yang et alyang2019exposing focuses on the inconsistency of head poses in videos, comparing the differences between head poses estimated using all facial landmarks and those estimated using only the landmarks in the central region. Peng et alpeng2024deepfakes focuse on inter-frame gaze angles, obtaining gaze characteristics of each video frame and using a spatio-temporal feature aggregator to combine temporal gaze features, spatial attribute features, and spatial texture features as the basis for detection and classification.
$\bullet$ Inter-Frame Inconsistency. Methods zheng2021exploring ; gu2022delving ; yin2023dynamic ; yang2023masked ; choi2024exploiting ; xu2024towards based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et algu2022delving focuse on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et alyin2023dynamic design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et alyang2023masked approach forgery detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames. Choi et alchoi2024exploiting discover that the style variables in each frame of Deepfake work change. Based on this, they developed a style attention module to focus on the inconsistency of the style latent variables between frames.
$\bullet$ Multimodal Inconsistency. The core idea behind multimodal detection algorithms is to make judgments based on the flow of prior information from multiple attributes rather than solely considering the image or audio differences of individual characteristics in each frame. The consideration of audio-visual modal inconsistency has received extensive research in various methods haliassos2022leveraging ; cozzolino2023audio ; feng2023self ; yang2023avoid . POI-Forensics cozzolino2023audio proposes a deep forgery detection method based on audio-visual authentication, utilizing contrastive learning to learn the most distinctive embeddings for each identity in moving facial and audio segments. AVoiD-DF yang2023avoid embeds spatiotemporal information in a spatiotemporal encoder and employs a multimodal joint decoder to fuse multimodal features and learn their inherent relationships. Subsequently, a cross-modal classifier is applied to detect disharmonious operations within and between modalities. Agarwal et alagarwal2021detecting describe a forensic technique for detecting fake faces using static and dynamic auditory ear characteristics. Indeed, multimodal detection methods are currently a hotspot in forgery detection research.
3.2.3 Frequency Domain
Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works qian2020thinking ; frank2020leveraging ; li2021frequency ; miao2022hierarchical ; tan2024frequency utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, F3-Net qian2020thinking proposes a dual-branch framework. One frequency-aware branch utilizes Frequency-aware Image Decomposition (FAD) to learn subtle forgery patterns in suspicious images. In contrast, the other branch aims to extract high-level semantics from Local Frequency Statistics (LFS) to describe the frequency-aware statistical differences between real and forged faces. HFI-Net miao2022hierarchical consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Tan et altan2024frequency introduce a novel frequency-aware approach called FreqNet, which focuses on the high-frequency information of images and combines it with a frequency-domain learning module to learn source-independent features. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration masi2020two ; guo2023constructing , and Guo et alguo2023constructing design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
3.2.4 Data Driven
Data-driven forgery detection focuses on learning specific patterns and features from extensive image or video datasets to distinguish between genuine and potentially manipulated images. Some methods hsu2018learning ; yu2019attributing believe that images generated by specific models possess unique model fingerprints. Based on this belief, forgery detection can be achieved by focusing on the model’s training. In addition, FakeSpotter wang2021fakespotter introduces the Neuron Coverage Criterion to capture layer-wise neuron activation behavior. It monitors the neural behavior of a deep face recognition system through a binary classifier to detect fake faces. There are also methods zhao2021learning ; huang2023implicit that attempt to classify the sources of different components in an image. For instance, Huang et alhuang2023implicit think that the difference between explicit and implicit identity helps detect face swapping. There are numerous data-driven methods dang2020detection ; hu2022finfer ; guo2023hierarchical ; guo2023controllable , and it is not feasible to discuss each one in detail here.
3.3 Specific Related Domains
In this section, we briefly review related popular tasks beyond the deepfake generation, such as Face Super-resolution (Sec. 3.3.1), Portrait Style Transfer (Sec. 3.3.2), Body Animation (Sec. 3.3.3), and Makeup Transfer (Sec. 3.3.4).
3.3.1 Face Super-resolution
$\bullet$ Convolutional Neural Networks. Early works chen2019sequential ; hu2019face ; lu2020global on facial super-resolution based on CNNs aims to leverage the powerful representational capabilities of CNNs to learn the mapping relationship between low-resolution and high-resolution images from training samples. Depending on whether they focus on local details of the image, they can be divided into global methods chen2019sequential ; huang2017wavelet , local methods hu2019face , and mixed methods lu2020global .
$\bullet$ Generative Adversarial Network. GAN aims to achieve the optimal output result through an adversarial process between the generator and the discriminator. This type of method zhang2019ranksrgan ; pan2024lpsrgan currently dominates the field for flexible and efficient architecture.
3.3.2 Portrait Style Transfer
$\bullet$ Generative Adversarial Network. The most mature style transfer algorithm is the GAN-based approach abdal20233davatargan ; xu2022your3demoji ; jiang2023scenimefy . However, due to the relatively poor stability of GANs, it is common for the generated images to contain artifacts and unreasonable components. 3DAvatarGAN abdal20233davatargan bridges the pre-trained 3D-GAN in the source domain with the 2D-GAN trained on an artistic dataset to achieve cross-domain generation. Scenimefy jiang2023scenimefy utilizes semantic constraints provided by text models like CLIP to guide StyleGAN generation and applies patch-based contrastive style loss to enhance stylization and fine details further.
$\bullet$ Diffusion-based methods liu2023portrait ; hur2024expanding ; lei2023diffusiongan3d represent the generative process of cross-domain image transfer using diffusion processes. DiffusionGAN3D lei2023diffusiongan3d combines 3DGAN chan2022efficient with a diffusion model from text to graphics, introducing relative distance loss and learnable tri-planes for specific scenarios to further enhance cross-domain transformation accuracy.
3.3.3 Body Animation
$\bullet$ Generative Adversarial Network. GAN-based approaches zhao2022thin ; jain2023vgflow ; zhou2022cross aim to train a model to generate images whose conditional distribution resembles the target domain, thus transferring information from reference images to target images. CASD zhou2022cross is based on a style distribution module using a cross-attention mechanism, facilitating pose transfer between source semantic styles and target poses. VGFlow jain2023vgflow introduces a visibility-guided flow module to preserve texture and perform style manipulation concurrently. However, existing methods still rely considerably on training samples, and exhibit decreased performance when dealing with actions in rare poses.
$\bullet$ Diffusion-based. The task of body animation using diffusion models aims to utilize diffusion processes to generate the propagation and interaction of movements between body parts based on a reference source. This approach wang2023leo ; ma2023follow ; hu2023animate represents a current hot topic in research and implementation. LEO wang2023leo focuses on the spatiotemporal continuity between generated actions, employing the Latent Motion Diffusion Model to represent motion as a series of flow graphs during the generation process. Animate Anyone hu2023animate harnesses the powerful generation capabilities of stable diffusion models combined with attention mechanisms to generate high-quality character animation video.
3.3.4 Makeup Transfer
$\bullet$ Graphics-based Approaches. Before the introduction of neural networks, traditional computer graphics methods guo2009digital ; li2015simulating use image gradient editing and physics-based operations to understand the semantics of makeup. By decomposing the input image into multiple layers, each representing different facial information, traditional methods would distort the reference facial image onto the non-makeup target using facial landmarks for each layer. However, due to the limitations of manually designed operators, the output images from traditional methods often appear unnatural, with noticeable artifacts. Additionally, there is a tendency for background information to be modified to some extent.
$\bullet$ Generative Adversarial Network. Early deep learning-based methods gu2019ladn aim at fully automatic makeup transfer. However, these methods chang2018pairedcyclegan ; chen2019beautyglow exhibit poor performance when faced with significant differences in pose and expression between the source and target faces and are unable to handle extreme makeup scenarios well. Some methods jiang2020psgan ; liu2021psgan++ ; zhong2023sara proposes their solutions, PSGAN++ liu2021psgan++ comprises the Makeup Distillation Network, Attentive Makeup Morphing module, Style Transfer Network, and Identity Extraction Network, further enhancing the ability of PSGAN jiang2020psgan to perform targeted makeup transfer with detail preservation. ELeGANt yang2022elegant , CUMTGAN hao2022cumtgan , and HT-ASE li2023hybrid explore the preservation of detailed information. ELeGANt yang2022elegant encodes facial attributes into pyramid feature maps to retain high-frequency information. Matching facial semantic information, which involves rendering makeup styles onto semantically corresponding positions of a target image, is often overlooked. SSAT sun2022ssat introduces the SSCFT module and weakly supervised semantic loss for accurate semantic correspondence. SSAT++ sun2023ssat further improves color fidelity matching, but both models are complex with high training complexity. In addition, BeautyREC yan2023beautyrec , based on the Transformer with long-range visual dependencies, achieves efficient global makeup transfer with significantly reduced overall model parameters compared to previous works.
4 Benchmark Results
We firstly introduce the evaluation metrics commonly used for each deepfake tasks, and then we evaluate the performance of representative methods for each reviewed field on the most widely used datasets with the data sourced from the original respective papers. Considering the differences in training datasets, testing datasets, and metrics used by different approaches, we strive to compare them fairly in each table to the greatest extent.
4.1 Metrics
$\bullet$ Face Swapping. The most commonly used objective evaluation metrics for face swapping include ID Ret, Expression Error, Pose Error, and FID. ID Ret is calculated by a pre-trained face recognition model wang2018cosface , measuring the Euclidean distance between the generated face and the source face. A higher ID Ret indicates better preservation of identity information. Expression and pose errors quantify the differences in expression and pose between the generated face and the source face. These metrics are evaluated using a pose estimator ruiz2018fine and a 3D facial model deng2019accurate , extracting expression and pose vectors for the generated and source faces. Lower values for expression error and pose error indicate higher facial expression and pose similarity between the swapped face and the source face. FID  heusel2017gans is used to assess image quality, with lower FID values indicating that the generated images closely resemble authentic facial images in appearance.
$\bullet$ Face Reenactment. Face reenactment commonly uses consistent evaluation metrics, including CSIM, SSIM wang2004image , PSNR, LPIPS zhang2018unreasonable , LMD chen2018lip , and FID. CSIM describes the cosine similarity between the generated and source faces, calculated by ArcFace deng2019arcface , with higher values indicating better performance. SSIM, PSNR, LPIPS, and FID are used to measure the quality of synthesized images. SSIM measures the structural similarity between two images, with higher values indicating a closer resemblance to natural images. PSNR quantifies the ratio between a signal’s maximum possible power and noise’s power, indicating higher quality for higher values. LPIPS assesses reconstruction fidelity using a pre-trained AlexNet krizhevsky2012imagenet to extract feature maps for similarity score computation. As mentioned earlier, FID is used to evaluate image quality. LMD assesses the accuracy of lip shape in generated images or videos, with lower values indicating better model performance.
$\bullet$ Talking Face Generation. Expanding upon face reenactment metrics, talking face generation incorporates additional metrics, including M/F-LMD, Sync, LSE-C, and LSE-D. The LSE-C and LSE-D are usually used to measure lip synchronization effectiveness prajwal2020lip . The landmark distances on the mouth (M-LMD) chen2019hierarchical and the confidence score of SyncNet (Sync) measure synchronization between the generated lip motion and the input audio. F-LMD computes the difference in the average distance of all landmarks between predictions and ground truth (GT) as a measure to assess the generated expression. In addition, there are some meaningful metrics, such as LSE-C and LSE-D for measuring lip synchronization effectiveness prajwal2020lip , AVD li2023one for evaluating identity preservation performance, AUCON ha2020marionette for assessing facial pose and expression jointly, and AGD doukas2023free for evaluating eye gaze changes. These newly proposed evaluation metrics enrich the performance assessment system by targeting various aspects of the model’s performance.
$\bullet$ Facial Attribute Editing. The standard evaluation metrics used in face attribute manipulation are FID, LPIPS zhang2018unreasonable , KID binkowski2018demystifying , PSNR and SSIM. KID is one of the image quality assessment metrics commonly used in face editing work and other generative modeling tasks to quantify the difference in distribution between the generated image and the actual image, with lower KID values indicating better model performance. Some text-guide work will also use the CLIP Score to measure the consistency between the output image and the text, calculated as the cosine similarity between the normalized image and the text embedding. Higher values of CLIP Score indicate better consistency of the generated image with the corresponding text sentence.
4.2 Main Results on Deepfake Generation
We select widely used metrics for each task and tabulate the performance for representative methods.
$\bullet$ Results on Face Swapping. Table 7 displays the performance evaluation results of some representative models on the Face Swapping task using the FF++ rossler2019faceforensics++ dataset. WSC-Swapren2023reinforced captures facial region external attribute information and internal identity information through two independent encoders, which achieves good identity preservation and facial pose retention. However, it exhibits sub-optimal performance in facial expression error metrics.
$\bullet$ Results on Face Reenactment. Table 8 and Table 9 show the performance evaluation results on the VoxCeleb nagrani2017voxceleb dataset for self-reenactment and cross-subject reenactment, respectively. AVFR-GAN agarwal2023audio achieves better performance by using the multimodal modeling. 10 presents quality assessment results on the VoxCeleb2 dataset. HiDe-NeRF li2023one represents 3D scenes using canonical appearance fields and implicit deformation fields. It achieves accurate facial attribute modeling by explicitly decoupling facial pose and expression attributes using the deformation module.
Table 7: Results of representative face swapping methods on FF++. Notations: ➊ CelebA-HQ, ➋ FFHQ, ➌ VGGFace, ➍ VGGFace2, ➎ VoxCeleb2.
Methods Train Test: FF++
ID Ret.(%)↑ Exp Err.↓ Pose Err.↓ FID↓
FaceShifter li2019faceshifter ➊➋➌ 97.38 2.06 2.96 -
SimSwap chen2020simswap 92.83 - 1.53 -
FaceInpainter li2021faceinpainter ➊➋➌ 97.63 - 2.21 -
HifiFace wang2021hififace 98.48 - 2.63 -
RAFSwap xu2022region 96.70 2.92 2.53 -
Xu et alxu2022high 90.05 2.79 2.46 -
DiffSwap zhao2023diffswap 98.54 5.35 2.45 2.16
FlowFace zeng2023flowface ➊➋➍ 99.26 - 2.66 -
FlowFace++ zhang2023flowface++ ➊➋➍ 99.51 - 2.20 -
StyleIPSB jiang2023styleipsb 95.05 2.23 3.58 -
StyleSwap xu2022styleswap ➌➎ 97.05 5.28 1.56 2.72
WSC-Swap ren2023reinforced ➊➋➌ 99.88 5.01 1.51 -
Table 8: Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 kellnhofer2019gaze360 , ➎ MPIIGaze zhang2015appearance , ➏ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
Methods Train Test: VoxCeleb
CSIM↑ PSNR↑ LPIPIS↓ FID↓ SSIM↑
HyperReenact bounareli2023hyperreenact 0.710 - 0.230 27.10 -
DG hsu2022dual 0.831 - - 22.10 0.761
AVFR-GAN agarwal2023audio - 32.20 - 08.48 0.824
Free-HeadGAN doukas2023free ➊➌➍➎ 0.810 22.16 0.100 35.40 -
HiDe-NeRF li2023one ➊➋➐ 0.931 21.90 0.084 - 0.862
Table 9: Results of representative face reenactment methods on VoxCeleb for the cross-identity reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 kellnhofer2019gaze360 , ➎ MPIIGaze zhang2015appearance , ➏ TalkingHead-1KH.
Methods Train Test : VoxCeleb
CSIM↑ AVD↓ AUCON↑ FID↓ AGD↓
HyperReenact bounareli2023hyperreenact 0.680 - - - -
AVFR-GAN agarwal2023audio - - - 09.05 -
Free-HeadGAN doukas2023free ➊➌➍➎ 0.789 - - 53.90 13.1
HiDe-NeRF li2023one ➊➋➏ 0.786 0.012 0.971 57.00 -
Table 10: Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRW, ➍ CelebV-HQ, ➎ TalkingHead-1KH.
Methods Train Test: VoxCeleb2
CSIM↑ PSNR↑ LMD↓ FID↓ SSIM↑
PC-AVS zhou2021pose ➋➌ - - 6.880 - 0.886
GC-AVT liang2022expressive - - 2.757 - 0.739
Wang et alwang2023memory - 28.92 1.830 - 0.830
PECHead gao2023high ➋➍➎ 1.590 - 23.05 -
DG hsu2022dual 0.721 - - 51.79 0.540
HiDe-NeRF li2023one ➊➎ 0.787 - - 61.00 -
$\bullet$ Results on Talking Face Generation. Table 11 displays the performance results of various talking face generation approaches on the MEAD wang2020mead dataset since 2023. AMIGO zhai2023talking achieves promising results among all methods, which utilizes seq2seq to generate facial landmarks for emotion tagging, matching mouth movements with emotional features. Additionally, it employs a landmark-to-image translation network to create facial images with fine textures.
$\bullet$ Results on Facial Attribute Editing.  Table 12 evaluates the quality level of generated images using FID, and Table 13 assesses facial reconstruction capabilities using PSNR, LPIPS, and SSIM. Due to different training and testing datasets, quantitative fair comparisons are not possible that just serve as performance demonstrations.
Table 11: Evaluation results of the models involved in talking face generation on the MEAD dataset. Notations: ➊ MEAD, ➋ LRW, ➌ VoxCeleb2, ➍ HDTF.
Method Train Test: MEAD
CSIM↑ LMD↓ M/F-LMD↓ Sync↑ FID↓ PSNR/SSIM↑
Xu et alxu2023high 0.83 2.36 - 3.500 15.91 30.09/0.850
EMMN tan2023emmn ➊➋ - - 2.780/2.870 3.570 - 29.38/0.660
AMIGO zhai2023talking ➊➌ - 2.44 2.140/2.440 - 19.59 30.29/0.820
SLIGO sheng2023stochastic 0.88 1.83 - 3.690 - 00000-/0.790
Gan et algan2023efficient ➊➌ - - 2.250/2.470 - 19.69 21.75/0.680
DreamTalk ma2023dreamtalk ➊➌➍ - - 2.910/1.930 3.780 - 00000-/0.860
SPACE gururani2023space ➊➌ - - - 3.610 11.68 0-
TalkCLIP ma2023talkclip - - 3.601/2.415 3.773 - 00000-/0.829
Table 12: FID evaluation of different methods. Notations: ➊ FFHQ, ➋ CelebA-HQ, ➌ MM CelebA-HQ, ➍ CelebA, ➎ CelebAText-HQ, ➏ Self-build.
Methods Type Train Test FID↓
FENeRF sun2022fenerf GANs+NeRF 12.10
FENeRF sun2022fenerf GANs+NeRF 28.20
AnyFace sun2022anyface GANs 56.75
AnyFace sun2022anyface GANs 50.56
TextFace hou2022textface GANs 22.81
TG-3DFace yu2023towards GANs 52.21
TG-3DFace yu2023towards GANs 39.02
HifaFace gao2021high GANs ➊➋ ➊➋ 04.04
GuidedStyle hou2022guidedstyle GANs 41.79
4.3 Main Results on Forgery Detection
The mainstream evaluation metrics for forgery detection technology are ACC and AUC. Table 14 presents the ACC and AUC metrics for some detection models trained on FF++ rossler2019faceforensics++ and tested on FF++ (HQ) and FF++ (LQ). LipForensics haliassos2021lips exhibits robust performance on the strongly compressed FF++ (LQ), while Guo et alguo2023controllable perform best on FF++ (HQ). Table 15 shows the cross-dataset evaluation using DFDC dolhansky2020deepfake , Celeb-DF li2020celeb , Celeb-DFv2 li2020celeb , and DeeperForensics-1.0 jiang2020deeperforensics as validation sets. AVoiD-DF yang2023avoid and Zhao et alzhao2021learning demonstrate excellent generalization ability, but there is still significant room for improvement in these datasets. However, overall, there is room for improvement in the evaluation performance of forgery detection models on the DFDC, which has a larger sample size and more complex forgery methods.
Table 13: Results of facial reconstruction capabilities in facial attribute editing work. Notations: ➊ FFHQ, ➋ VoxCeleb, ➌ VoxCeleb2, ➍ CelebA-HQ.
Methods Type Train Test PSNR↑ LPIPS↓ SSIM↑
Preechakul et alpreechakul2022diffusion Difussion - 0.0110 0.991
Kim et al. kim2023diffusion Difussion - 0.0450 0.922
FDNeRF zhang2022fdnerf GANs+NeRF - 0.1420 0.821
IA-FaceS huang2023ia GANs ➊➍ 22.34 0.2240 0.642
IA-FaceS huang2023ia GANs ➊➍ 22.43 0.0384 0.659
Table 14: Results of the self-dataset performance on FF++. Notations: HQ (Mild compression), LQ (Heavy compression).
Methods Train FF++ (LQ) FF++ (HQ)
ACC(%) AUC(%) ACC(%) AUC(%)
F3-Net qian2020thinking FF++ 93.02 95.80 98.95 99.30
Masi et almasi2020two FF++ 86.34 - 96.43 -
Zhao et alzhao2021multi FF++ 88.69 90.40 97.60 99.29
FDFL li2021frequency FF++ 89.00 92.40 96.69 99.30
LipForensics haliassos2021lips FF++ 94.20 98.10 98.80 99.70
RECCE cao2022end FF++ 91.03 95.02 97.06 99.32
Guo et alguo2023controllable FF++ 92.76 96.85 99.24 99.75
MRL yang2023masked FF++ 91.81 96.18 93.82 98.27
Table 15: Results of cross-dataset performance evaluation on four datasets DFDC, Celeb-DF (CDF), Celeb-DFv2 (CDFv2), and DeeperForensics-1.0 (DFo). Evaluation indicator is AUC. Notations: ➊ FF++, ➋ FF++(Real), ➌ FF++(HQ), ➍ FF++(LQ), ➎ Self-build, ➏ SR-DF wang2022m2tr , ➐ DFDCp, ➑ FakeAVCeleb, ➒ DefakeAVMiT.
Methods Train DFDC CDF CDFv2 DFo
Face X-ray li2020face ➊➎ 80.92 80.58 - -
Face X-ray li2020face 71.15 74.76 - -
Masi et almasi2020two - 73.41 - -
Zhao et alzhao2021multi 67.44 - - -
Zhao et alzhao2021learning 67.52 98.30 90.03 99.41
Zheng et alzheng2021exploring 74.00 - 86.90 98.80
LipForensics haliassos2021lips 73.50 - 82.40 97.60
M2TR wang2022m2tr - 82.10 - -
M2TR wang2022m2tr - 68.20 - -
RECCE cao2022end 69.06 68.71 - -
SBIs shiohara2022detecting - - 90.79 -
SBIs shiohara2022detecting 72.42 - 93.18 -
RealForensics haliassos2022leveraging 75.90 - 86.90 99.30
Guo et alguo2023controllable 81.65 84.97 - -
Yin et alyin2023dynamic 73.08 71.36 - -
MRL yang2023masked 71.53 83.58 - -
AVoiD-DF yang2023avoid 80.60 - - -
AVoiD-DF yang2023avoid 90.30 - - -
5 Future Prospects
This section provides a short discussion of future research directions that can facilitate and envision the development of deepfake generation and detection.
$\bullet$ Face Swapping. Generalization is a significant issue in face swapping models. While many models demonstrate excellent performance on their training sets, there is often noticeable performance degradation when applied to different datasets during testing. In addition, beyond the common evaluation metrics, various face swapping works employ different evaluation metric systems, lacking a unified evaluation protocol. This absence hinders researchers from intuitively assessing model performance. Therefore, establishing comprehensive experimental and evaluation frameworks is crucial for fair comparisons, and driving progress in the field.
$\bullet$ Face Reenactment. Existing methods have room for improvement, facing three main challenges: convenience, authenticity, and security. Many approaches struggle to balance lightweight deployment and generating high-quality reenactment effects, hindering the widespread adoption of facial reenactment technology in industries. Moreover, several methods claim to achieve high-quality facial reenactment, but they exhibit visible degradation in output during rapid pose changes or extreme lighting conditions in driving videos. Additionally, the computational complexity, consuming significant time and system resources, poses substantial challenges for practical applications.
$\bullet$ Talking Face Generation. Current methods strive to enhance the realism of generated conversational videos. However, they lack fine-grained control over the emotional nuances of the conversation, where the matching of emotional intonation to audio and semantic content is not precise enough, and the control over emotional intensity is too coarse. In addition, the realistic correlation between head pose and facial expression movement seems insufficiently considered. Lastly, for text or audio with intense emotions, noticeable artifacts in head movement still occur to a significant extent.
$\bullet$ Facial Attribute Editing. Currently, mainstream facial attribute editing employs the decoupling concept based on GANs and diffusion models are gradually being introduced into this field. The primary challenge is effectively separating the facial attributes to prevent unintended processing of other facial features during attribute editing. Additionally, there needs to be a universally accepted benchmark dataset and evaluation framework for fair assessments of facial editing.
$\bullet$ Forgery Detection. With the rapid development of facial forgery techniques, the central challenge in face forgery detection technology is accurately identifying various forgery methods using a single detection model. Simultaneously, ensuring that the model exhibits robustness when detecting forgeries in the presence of disturbances such as compression is crucial. Most detection models follow a generic approach targeting common operational steps of a specific forgery method, such as the integration phase in face swapping or assessing temporal inconsistencies, but this manner limits the model’s generalization capabilities. Moreover, as forgery techniques evolve, forged videos may evade detection by introducing interference during the detection process.
6 Conclusion
This survey comprehensively reviews the latest developments in the field of deepfake generation and detection, which is the first to cover a variety of related fields thoroughly and discusses the latest technologies such as diffusion. Specifically, this paper covers an overview of basic background knowledge, including concepts of research tasks, the development of generative models and neural networks, and other information from closely related fields. Subsequently, we summarize the technical approaches adopted by different methods in the mainstream four generation and one detection fields, and classify and discuss the methods from a technical perspective. In addition, we strive to fairly organize and benchmark the representative methods in each field. Finally, we summarize the current challenges and future research directions for each field.
Acknowledgement. This work is supported by the National Natural Science Foundation of China (No. 62371189).